# Efficient Prior Selection in Gaussian Process Bandits with Thompson Sampling

### v2

@@ Appendix C Additional experimental details
For all the real-world datasets, sensors containing any null measurements have been filtered out.
@@
The PNW precipication data consists of daily precipication data from 1949 to 1950 across 167 $50\times 50$ km regions in the Pacific Northwest. The goal is to find the region with the highest precipitation for any given day. The training data consists of the measurements made prior to 1980 and the test data consists of the measurements between 1980 and 1994. The original data is stated to be given in mm/day however the data seems to be off by a factor of 10. We rescale the data to a log-scale using $\log(\cdot/10+0.1)$, similar to Krause et al. (2008). The noise variance is set to $\sigma^{2}=0.41^{2}$.
In the Intel experiment, we removed one outlier seed. All methods had a final cumulative regret around 6000°C, note that the average for the worst performing model across the other seeds was $\approx 150$°C. The outlier is shown Fig.˜8. We can see that one of the sensors display very high temperatures compared to all other sensors, which is why all methods performed poorly on this seed. It should be noted that many of the sensors in the Intel data logged degrees above 100°C after a certain time - likely due to sensor failure rather boiling temperatures in an office environment. Also note that these days were excluded from both our training and test data. The outlier could be an indication that this particular sensor was starting to fail earlier than others.
@@
Appendix D Additional experimental results
@@
First, we include the mean number of priors in $P_{t}$ for all experiments in Fig.˜9. Similarly, we include the average entropy of the hyperposterior for all experiments in Fig.˜10. For the lengthscale, subspace, PeMS and PNW precipication experiments, hardly any priors are eliminated. In contrast, the hyperposterior entropy concentrates rapidly across all experiments with the subspace and PNW precipication having the most and least concentrated hyperposterior.
We include the full set of confusion matrices for the lengthscale and subspace experiments in Fig.˜11. In the lengthscale experiments, we observe that PE-GP-UCB and -TS oversample the shortest lengthscale. This is similar to the kernel experiment where the Matérn 3/2 kernel was also oversampled. However, we see that HP-GP-TS and MAP GP-TS do not suffer from this optimistic bias. In the subspace experiment, HP- and MAP GP-TS have an accuracy of around 96% where as PE-GP-TS and -UCB have accuracies 30% and 36% respectively. The priors are equivalent up to coordinate permutations and are therefore difficult. The PE-methods do not oversample any specific prior but commit to much time to exploring along the irrelevant dimensions.
In Tables˜1 and 2, the total regret for the lengthscale and subspace scaling experiments are shown.
In Fig.˜12, we visualize the median cumulative regret on the real-world data experiments. Similarly, in Fig.˜13, we show further quantiles of the final cumulative regret. MAP and HP-GP-TS exhibit similar median regret across the experiments but MAP GP-TS has a larger variance and tail values.
@@
Table 1: Average total regret and $\pm 1$ standard error for the lengthscale experiment as $|P|$ increases.
@@
Table 2: Average total regret and $\pm 1$ standard error for the subspace experiment as $|P|$ increases.
@@
Figure 12: Median cumulative regret on Intel temperature data (top), PeMS speed data (middle) and PNW precipication data (bottom). Errorbars correspond to first and last decile.
(a) Intel
(b) PeMS
(c) PNW Precipication
Figure 13: Quantiles of the final cumulative regret on the real-world data experiments. The median is shown with a black line. The whiskers correspond to the 5th and 95th percentile and the lower and upper edges of the boxes show the first and third quantile.

### v3

@@ Appendix E Additional experimental details
For the kernel experiment, all kernels use a lengthscale of $1.0$ and are scaled s.t. $k(x,\tilde{x})\leq 1$. In addition, the mean function for all priors is zero everywhere. For the subspace experiment, the total dimensions $d=16$ but each prior $p_{i}$ assumes $f(x)$ depends on $d_{s}=4$ subdimensions: $[i,i+1,i+2,i+3]$ for $i\in[5]$. Dimensions larger than $5$ are wrapped around 1, i.e. $((j-1)\mod 5)+1$, such that the priors are equally difficult to distinguish and optimize. The prior elimination methods use $\delta=0.05$ across all experiments, including the oracle methods. During every iteration $t$, SCoreBO samples $M$ priors from the hyperposterior $P_{t}$ and samples $N$ optimizers $x^{*},f^{*}$ for each prior sampled through posterior sampling. We use $M=16$ and $N=12$ for SCoreBO in all experiments. We use the implementation of the SCoreBO acquisition function in BoTorch (Community) (Balandat et al., 2020). To make the implementation fast with GPUs, we set linear_operator.settings.stable_qr_cpu_threshold to 8 in order to avoid QR-factorization being performed on CPU (Gardner et al., 2018; Pleiss et al., 2022, 2025). To avoid out of memory issues, we replace the default torch.matmul in DefaultPredictionStrategy._exact_predictive_covar_inv_quad_form_root (from gpytorch.models.exact_prediction_strategies) with an equivalent torch.einsum (Gardner et al., 2018). Since the priors in our experiments are discrete, we compute the hyperposterior exactly and sample from it directly.
For all the real-world datasets, sensors containing any null measurements are filtered out.
@@
The PNW precipitation data consists of daily precipitation data from 1949 to 1994 across 167 $50\times 50$ km regions in the Pacific Northwest. The goal is to find the region with the highest precipitation for any given day. The training data consists of the measurements made prior to 1980 and the test data consists of the measurements between 1980 and 1994. The original data is stated to be given in mm/day however the data seems to be off by a factor of 10. We rescale the data to a log-scale using $\log(\cdot/10+0.1)$, similar to Krause et al. (2008). The noise variance is set to $\sigma^{2}=0.41^{2}$.
In the Intel experiment, we removed one outlier seed. All methods had a final cumulative regret around 6000°C on this instance, note that the average for the worst performing model across the other seeds was $\approx 150$°C. The outlier is shown Figure˜9. We can see that one of the sensors display very high temperatures compared to all other sensors, which is why all methods performed poorly on this seed. It should be noted that many of the sensors in the Intel data logged degrees above 100°C after a certain time - likely due to sensor failure rather boiling temperatures in an office environment. Also note that these days were excluded from both our training and test data. The outlier could be an indication that this particular sensor was starting to fail earlier than others.
@@
Appendix F Additional experimental results
@@
Table 1: Average total regret and $\pm$ 1 standard error for the synthetic and real-world data experiments. The algorithms with the lowest regret (excluding oracle algorithms) are highlighted in bold.
Algorithm Synthetic Real-world data
Kernel Lengthscale Subspace Intel PeMS PNW Precip.
MAP GP-TS 84.3 $\pm$ 8.4 30.2 $\pm$ 1.2 87.2 $\pm$ 1.0 73.8 $\pm$ 7.7 1635.0 $\pm$ 129.3 178.4 $\pm$ 6.9
HP-GP-TS 39.2 $\pm$ 1.4 31.4 $\pm$ 1.0 88.3 $\pm$ 0.9 54.1 $\pm$ 3.0 1327.8 $\pm$ 107.9 167.7 $\pm$ 5.2
PE-GP-TS 62.0 $\pm$ 0.6 61.8 $\pm$ 0.5 177.1 $\pm$ 1.4 106.5 $\pm$ 2.1 1214.2 $\pm$ 81.5 200.9 $\pm$ 4.0
PE-GP-UCB 121.6 $\pm$ 1.2 114.2 $\pm$ 0.6 389.0 $\pm$ 1.5 173.0 $\pm$ 2.7 2159.2 $\pm$ 48.4 506.2 $\pm$ 2.6
Oracle GP-TS 35.0 $\pm$ 1.1 28.1 $\pm$ 0.8 86.0 $\pm$ 1.0
Oracle GP-UCB 68.5 $\pm$ 1.9 48.3 $\pm$ 1.2 217.3 $\pm$ 1.0
SCoreBO 180.4 $\pm$ 7.7 180.8 $\pm$ 5.8 106.6 $\pm$ 0.9 256.8 $\pm$ 9.3 3460.2 $\pm$ 163.7 861.3 $\pm$ 21.0
EEI 39.0 $\pm$ 2.6 30.1 $\pm$ 2.1 88.3 $\pm$ 4.2 51.6 $\pm$ 4.8 1664.1 $\pm$ 137.7 196.5 $\pm$ 12.3
First, we provide the average total regret for the synthetic and real-world data experiments in Table˜1. We observe that HP-GP-TS either has the lowest regret or is within 1 standard error of the algorithm with the lowest regret across all the experiments. In Figure˜3, it can be noted had low regret early in the synthetic experiment but HP-GP-TS either catches up or almost catches up later in the experiments. We compare both algorithms with an extended time horizon of $T=1500$, the results are shown in Figures˜10 and 2. With the extended time horizon, HP-GP-TS achieves the lowest regret across all synthetic experiments. Although, EEI is still within 1 standard error on the lengthscale experiment.
Figure 10: Cumulative regret for synthetic experiments extended time horizon $T=1500$ with varying kernel (left), lengthscale (center) and mean function (right). Errorbars correspond to $\pm 1$ standard error.
Table 2: Average total regret and $\pm$ 1 standard error for the synthetic experiments with longer horizon $T=1500$. The algorithms with the lowest regret (excluding oracle algorithms) are highlighted in bold.
Algorithm Kernel Lengthscale Subspace
HP-GP-TS 49.1 $\pm$ 1.6 39.7 $\pm$ 1.2 103.4 $\pm$ 1.3
EEI 62.6 $\pm$ 6.2 45.7 $\pm$ 5.0 138.9 $\pm$ 9.2
@@
Next, we include the mean number of priors in $P_{t}$ for all experiments in Figure˜11. Similarly, we include the average entropy of the hyperposterior for all experiments in Figure˜12. For the lengthscale, subspace, PeMS and PNW precipitation experiments, hardly any priors are eliminated. In contrast, the hyperposterior entropy concentrates rapidly across all experiments with the subspace and PNW precipitation having the most and least concentrated hyperposteriors.
The cost of learning the prior, ${\sum_{t\in[T]}}{\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$, for HP-GP-TS is shown in Figure˜13. Whilst this term has not been bounded theoretically in Theorem˜4.2, we observe that empirically its growth diminishes much quicker than the cumulative regret. Since ${\sum_{t\in[T]}}{\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$ is one term in the upper bound for the regret of HP-GP-TS, learning the prior does not seem to increase the growth rate of the regret empirically.
Figure 13: The cost of learning the prior, ${\sum_{t\in[T]}}U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})$, for HP-GP-TS in the synthetic experiments. The solid blue line corresponds to the mean and the dashed lines corresponds to the first and last decile at each time step.
We include the full set of confusion matrices for the lengthscale and subspace experiments in Figure˜14. In the lengthscale experiments, we observe that PE-GP-UCB and -TS oversample the shortest lengthscale. This is similar to the kernel experiment where the Matérn 3/2 kernel was also oversampled. However, we see that HP-GP-TS and MAP GP-TS do not suffer from this optimistic bias. In the subspace experiment, HP- and MAP GP-TS have an accuracy of around 96% where as PE-GP-TS and -UCB have accuracies 30% and 36% respectively. Even though PE-GP-UCB has a higher accuracy than PE-GP-TS, it still has significantly higher regret. Additionally, the priors are equivalent up to coordinate permutations and therefore generate functions that are equally difficult to optimize. Unlike the kernel and lengthscale experiments, the PE-methods do not oversample any specific prior but commit too much time to exploring along the irrelevant dimensions.
In Tables˜3 and 4, the total regret for the lengthscale and subspace scaling experiments are shown.
@@
Table 3: Average total regret and $\pm 1$ standard error for the lengthscale experiment as $|P|$ increases. The algorithms with the lowest regret (excluding oracle algorithms) are highlighted in bold.
@@
SCoreBO 180.8 $\pm$ 5.8 240.3 $\pm$ 6.3 277.2 $\pm$ 6.8 281.6 $\pm$ 6.7 283.9 $\pm$ 6.8
EEI 30.1 $\pm$ 2.1 30.9 $\pm$ 2.2 29.4 $\pm$ 2.2 30.8 $\pm$ 2.6 32.1 $\pm$ 2.8
Table 4: Average total regret and $\pm 1$ standard error for the subspace experiment as $|P|$ increases. The algorithms with the lowest regret (excluding oracle algorithms) are highlighted in bold.
@@
SCoreBO 106.6 $\pm$ 0.9 108.2 $\pm$ 0.8 108.9 $\pm$ 0.7 109.5 $\pm$ 0.7
EEI 88.3 $\pm$ 4.2 81.3 $\pm$ 3.8 82.5 $\pm$ 3.6 81.3 $\pm$ 3.8